Ammonia is an important compound and useful to industries such as

agricultural, pharmaceuticals, textile, plastics, petroleum, mining, and commercial. In the ammonia production, Haber process is utilized and it is shown as,

N2(g) + 3H2(g) → 2NH3(g).


How many grams of hydrogen gas are needed to produce 10 000.0 g of ammonia?



1
Expert's answer
2021-11-02T02:18:07-0400

N2 + 3H2 → 2NH3

Calculate moles of NH3 in 10.00 g.

n = m/M, where;

n = moles = ?

m = mass(g) = 10.00 g NH3

M = molar mass(g/mol) = (1 × 14.007 g/mol N) + (3 × 1.008 g/mol H) = 17.031 g/mol NH3

n NH3 = 10.00 g NH3/17.031 g NH3/mol NH3 = 0.58716 mol NH3

Since molar mass is a fraction (g/mol), you can divide by multiplying by its reciprocal (mol/g). This makes it easier to see how grams cancel, leaving moles.

n NH3 = 10.00 g NH3 × 1 mol NH3/17.031 g NH3 = 0.58716 mol NH3

Use stoichiometry to calculate the mole H2 required to produce 0.58716 mol NH3.

Multiply 0.58716 mol NH3 by the mole ratio between H2 and NH3 in the balanced equation, so that mol NH3 cancel, leaving mol H2.

0.58716 mol NH3 × 3 mol H2/2 mol N2 = 0.88074 mol H2

Calculate the mass of 0.88074 mol H2.

m = n × M

M H2 = (2 × 1.008 g/mol H) = 2.016 g/mol

m H2 = 0.88074 mol H2 × 2.016 g H2/mol H2 = 1.776 g H2 to four significant figures

1.776 g H2 is required to produce 10.00 g NH3, if N2 is in excess.